# Optical and UV Flares from Binary Black Hole Mergers in Active Galactic Nuclei

### v1

@@ Abstract
Stellar mass, binary black hole (BBH) mergers dominates the sources of gravitational wave (GW) events so far detected the LIGO/Virgo/KAGRA (LVK) experiment. The origin of these BBHs is unknown, and no electromagnetic (EM) counterpart has been undoubtedly associated to any of such GW events. The thin discs of active galactic nuclei (AGNs) might be viable environments where BBHs can form and at the same time produce observable radiation feedback. This paper presents new physically motivated light-curve (LC) solutions for thermal flares driven by the remnant of a BBH merger within the disc of an AGN. Following previous analyses, we consider that the BBH likely creates an under-density cavity in the disc prior to its coalescence. Depending on the merger conditions, the black hole (BH) remnant can leave the cavity, interact with the unperturbed disc, and drive a transient BH wind. The wind expels disc material that expands above and below the disc plane and we consider the emission of these plasma ejections as the EM counterpart to the merger GW. We model the LC of such eruptions as mini supernovae explosions finding that stellar mass merger remnants can drive distinguishable flares in optical and UV bands, with time lags of $\sim 10-40$ ($100-400$) days after the GW event, lasting for a few days (weeks) when occurring in AGN with central engines of $\sim 10^{6}$ ($10^{7}$) M. The Vera C. Rubin Observatory can potentially detect the optical component of these EM counterparts up to redshifts of $z\sim 0.5-1$, accordingly. The present LC models are applicable to localise flaring AGNs as sources of multi-messenger emission following GW alerts.
@@ I Introduction
Electromagnetic flares accompanying BBH mergers could unveil their astrophysical hosts. The search for such multi-messenger (MM) associations is the subject of current active research involving follow-up campaigns [8, 9, 10, 11, 12], statistical analyses of observed data [11, 13, 14, 15, 16, 17], as well as theoretical efforts to understand the production of detectable electromagnetic counterparts [18, 19, 20, 21, 22, 23, 24, 25]. GWs and radiation from a common origin provide the means of measuring the distance and redshift of the source independently. In this regard, MM events triggered by BBH mergers can also find cosmological applications [26, 27, 28, 29]. Contrary to mergers of compact objects involving neutron stars, BBH mergers do not produce radiation by themselves and require occurring in gas-rich environments to produce significant EM signals.
The central regions of active galactic nuclei (AGNs) possess suitable conditions to cluster stellar mass black holes [30, 31, 32] and to induce BBH pairing and merging [33, 34, 35, 36, 37]. The scenario of mergers in AGN discs appears to naturally explain puzzling features revealed by LVK measurements, such as the anti-correlation of mass ratio and effective spin [14], as well as the production of intermediate mass black holes as merger remnants [38, 33]. Interestingly, AGN thin discs are environments dense enough where BBH mergers can produce significant radiation feedback [34, 20, 22].
@@
Figure 1: Sketch of the EM counterpart scenario discussed in this work. Upper: Radial cut of an AGN thin disc where the BBH coalescence occurs within a pre-merger cavity. The merger remnant is kicked after coalescence and penetrates the unperturbed disc region. Middle: While travelling within the AGN disc, the remnant experiences a highly super-Eddington inflow of gas, leading to the production of a quasi-isotropic wind that expels the disc material above and below the disc plane. Lower: The plasma ejections produce an electromagnetic flare after expanding enough to allow the escape of internal photons.
This paper is organised as follows. In the next section, we derive an analytical model for the EM counterpart emission. In Section II.3, we discuss the model parameters of interest and present the spectra of the predicted emission. In Section IV, we derive optical light-curve profiles, compare them with the sensitivity of optical instruments. We present a summary and conclusions in Section V.
@@ II Electromagnetic counterpart: analytical considerations
Following such prediction and for analytical convenience, here we consider a BBH coalescence at the mid-plane of a thin disc of semiheight $h_{\mathrm{d}}$, and within a cylindrical cavity of radius $r_{\mathrm{cav}}$, as depicted in Figure 1.
@@
A burst of GWs is released at coalescence, and the merger remnant is born with a mass $\sim M_{\bullet}$ [^1] The mass of the remnant is typically slightly smaller than the mass of the BBH due to the release of gravitational waves at coalescence. Nevertheless, for practical purposes proposed in this work, we do not make a distinction between the masses of the BBH and the remnant. and with a likely gravitational kick velocity $v_{\mathrm{k}}$ in the frame of the coalescence. The initial condition of the coalescence is uncertain, thus we consider the kick direction as random.
Figure 2: Sketch of the geometry and the characteristic lengths analysed in this work. Upper: Radial cut of the AGN disc similar to the upper panel of Figure 1. The red and blue regions represent the volumes of the disc that we assume to be expelled above and below the disc plane. Lower: Edge-on view of the AGN disc at the coalescence site. The white circle represents the pre-merger cavity, and the dashed straight line represents the trajectory of the kicked remnant that penetrates the unperturbed disc. The parabolic-like area represents the pseudo-cometary structure discussed in the text, formed due to the interaction of the remnant-driven wind with the matter of the AGN disc.
@@
The magnitude of the kick velocity can range within $v_{\mathrm{k}}\sim$ [100, 2000] km s-1 [53, 54], which is generally highly supersonic with respect to the disc material. On the other hand, the radial drift velocity of the disc matter is highly subsonic [55], and thus the cavity crossing time of the remnant $r_{\mathrm{cav}}/v_{\mathrm{k}}$ is much shorter than the timescale of cavity replenishment [22]. Therefore, the interaction of the remnant with the unperturbed disc (and thus the possible production of electromagnetic feedback) crucially depends on the kick direction $\theta_{\mathrm{k}}$ with respect to the disc plane. We expect no interaction and hence no radiation counterpart for remnants with $\theta_{\mathrm{k}}\rightarrow\pi/2$. On the other hand, a significant interaction between the remnant and the unperturbed disc is expected for small $\theta_{\mathrm{k}}$ (see Figure 1), which is the case we focus on in this work.
@@
A kicked remnant of mass $M_{\bullet}$ penetrating the disc volume perturbs the disc material due to its gravitational influence. Such interaction can be initially assessed with Bondi-Hoyle-Lyttleton (BHL) theory. In a medium of gas density $\rho_{\mathrm{d}}$, the classical BHL accretion onto the kicked remnant is approximately $\dot{M}_{\mathrm{B}}\sim 4\pi\rho_{\mathrm{d}}(GM_{\bullet})^{2}/v_{\mathrm{k}}^{3}$. Normalising this inflow rate by the Eddington accretion limit of the remnant, gives for typical parameters of the problem:
@@
where $n_{\mathrm{d}}$ is the disc number density. Under such highly super-Eddington inflow, the remnant is expected to develop radiation-driven winds and/or bipolar collimated jets. Recent general relativistic magnetohydrodynamic (GRMHD) simulations of BHL accretion performed by [56] show the production of Blandford-Znajek jets [57] by spinning black holes (BHs) with a fixed spin direction, when traveling through a medium with plasma beta $\beta\gtrsim 10$. It is not clear, however, what the accretion disc’s $\beta$ is where BBHs likely coalesce. Furthermore, it is not clear whether the BBH coalescence produces a remnant with a fixed spin direction. BHs with precessing spin may launch precessing jets that could globally resemble wide-angle bipolar outflows. Given the aforementioned uncertainties, in the present work we consider that the remnant drives a quasi-spherical transient wind during its trajectory within the AGN disc. We assume that the remnant captures matter from the environment while simultaneously driving outflows. These inflow-outflow processes have been analysed in the context of dynamical friction of travelling black holes (BHs) with feedback [58, 59], and to explain the emission of ultra-luminous X-ray sources as supercritical accretors [60, 61].
@@
Once the remnant enters the disc, it takes a period to capture enough material to ignite the wind, and we estimate such period as the crossing time of the BHL radius, in the source frame $\Delta t^{\prime}_{\mathrm{HL}}=2GM_{\bullet}/v_{\mathrm{k}}^{3}$. Under this consideration, a necessary condition for the production of a remnant wind is:
$\frac{2GM_{\bullet}}{v_{\mathrm{k}}^{2}}<\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}},$ (3)
i.e., the remnant’s trajectory within the disc (RHS of Equation 3) must be larger than the remnant’s BHL radius (see Figure 2). In condition 3, $h_{\mathrm{d}}$ is the local disc semi-height, which we parameterise using the Shakura & Sunyaev (SS) disc model [55].
@@
The wind is quenched when the remnant exits the disc (i.e., when the black hole has insufficient material to fuel the outflow), leaving two expansive ejections of plasma above and below the accretion disc as depicted in Figure 1 (bottom panel). In this work, we consider the emission produced by such disc eruptions as the electromagnetic counterpart to the preceding merger gravitational waves (GWs), and we approximate the associated light curves (LCs) as those of explosive blast waves, similar to approaches in [62, 63, 24]. In Subsection II.1, we estimate the amount of matter and energy in the disc ejections, and in Subsection II.2 we assess their emission.
@@ II.1 The energy and mass of the disc ejections
During its passage through the disc, the remnant drives a transient outflow which we refer to as the ”BH wind”. This BH wind expels disc matter that emerges from the disc on the same and opposite sides where the remnant leaves the disc surface (see Figure 1, bottom). These ejections are referred to as the ”upper” and ”lower” eruptions, respectively.
@@
where $\rho_{\mathrm{d}}$ and $c_{\mathrm{s}}$ are the local density and speed of sound in the disc, which we obtain using the SS disc model.
@@
The total energy injected by the remnant is calculated as $E_{0}=L_{\mathrm{w}}\Delta t^{\prime}_{\mathrm{w}}$, where $\Delta t^{\prime}_{\mathrm{w}}$ is the time period of wind injection (source frame):
$\Delta t_{\mathrm{w}}=\frac{l_{\mathrm{d}}-l_{\mathrm{HL}}}{v_{\mathrm{k}}},$ (9)
where $l_{\mathrm{d}}=h_{\mathrm{d}}/\sin\theta_{\mathrm{k}}-r_{\mathrm{cav}}/\cos\theta_{\mathrm{k}}$ is the length of the BH trajectory within the disc, and $l_{\mathrm{HL}}=2GM_{\bullet}/v_{\mathrm{k}}^{2}$ is the distance travelled before the wind is ignited (see Figure 1). We assume equipartition of the total wind power for energising the upper and lower ejections of matter, which is a reasonable approximation for small kick angles $\theta_{\mathrm{k}}$. Based on the aforementioned considerations, the total energy of the plasma ejected above (+) and below (-) the disc plane is parameterised as:
$E_{0}=\eta_{\mathrm{w}}\eta_{\mathrm{in}}2\pi\rho_{\mathrm{d}}\frac{(GM_{\bullet}c)^{2}}{v_{\mathrm{k}}^{4}}\left[\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}}-\frac{2GM_{\bullet}}{v_{\mathrm{k}}^{2}}\right].$ (10)
We note that in the small $\theta_{\mathrm{k}}$ regime, lower limits for the wind injection period and for the energy stored in each plasma ejection (equations 9 and 10, respectively) can be estimated as (see Figure 2):
$\Delta t^{\prime}_{\mathrm{w,ll}}$ $\sim\frac{h_{\mathrm{d}}}{v_{\mathrm{k}}}\sim 5.721\,\,\mathrm{days}$
×$\times\left(\frac{M_{\mathrm{S}}}{10^{6}\mathrm{M}_{\odot}}\right)\left(\frac{a/R_{\mathrm{g}}}{10000}\right)\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)\left(\frac{v_{\mathrm{k}}}{300\mathrm{kms}^{-1}}\right)^{-1},$ (11)
and
$E_{\mathrm{0,ll}}$ $\sim\eta_{\mathrm{w}}\eta_{\mathrm{in}}2\pi\rho_{\mathrm{d}}h_{\mathrm{d}}\frac{(GM_{\bullet}c)^{2}}{v_{\mathrm{k}}^{4}}\sim 2.466\times 10^{47}\,\mathrm{erg}$
×$\times\left(\frac{\eta_{\mathrm{w}}}{0.005}\right)\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)\left(\frac{M_{\bullet}}{40\mathrm{M}_{\odot}}\right)^{2}\left(\frac{v_{\mathrm{k}}}{300\mathrm{kms}^{-1}}\right)^{-4}$
×$\times\left(\frac{M_{\mathrm{S}}}{10^{6}\mathrm{M}_{\odot}}\right)\left(\frac{a/R_{\mathrm{g}}}{10000}\right)\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)\left(\frac{n_{\mathrm{d}}}{10^{14}\mathrm{cm}^{-3}}\right),$ (12)
respectively.
The disc ejections are comprised of the material from the BH wind and swept-up disc material, as illustrated in Figure 1, upper. The wind material is accreted from the disc itself onto the remnant. Therefore, we estimate the masses of the upper and lower ejections as:
$M_{0}^{\pm}=\rho_{\mathrm{d}}V_{\mathrm{0}}^{\pm},$ (13)
where $\rho_{\mathrm{d}}$ is the density of the disc material surrounding the pre-merger cavity, and $V_{0}^{\pm}$ are the volumes in the disc above and below the trajectory of the remnant. The kick velocity $v_{\mathrm{k}}$ is typically highly supersonic with respect to the local sound speed of the disc, hence we expect the BH wind to form a bow shock structure. Considering the area $A_{\mathrm{b}}$ defined by the bow shock shape on a plane containing the BH trajectory, we obtain the upper volume $V_{0}^{+}$ as that of the ”wedge” with faces of areas $A_{\mathrm{b}}$ and $\cos\theta_{\mathrm{k}}A_{\mathrm{b}}$ (red region in Figure 1, middle). The lower volume $V_{0}^{-}$ is that of a prism with a base of $\cos\theta_{\mathrm{k}}A_{\mathrm{b}}$ plus the upper volume (blue region in Figure 1, middle):
$V_{0}^{\pm}=\cos\theta_{\mathrm{k}}A_{\mathrm{b}}\times\begin{cases}&\frac{1}{2}h_{\mathrm{d}}-\frac{1}{2}\tan\theta_{\mathrm{k}}r_{\mathrm{cav}}\;\;(\mbox{upper}),\\ \\ &\frac{3}{2}h_{\mathrm{d}}+\frac{1}{2}\tan\theta_{\mathrm{k}}r_{\mathrm{cav}}\;\;(\mbox{lower}).\end{cases}$ (14)
The bow shock area is calculated as:

### v2

@@ Abstract
Stellar mass, binary black hole (BBH) mergers dominates the sources of gravitational wave (GW) events so far detected by the LIGO/Virgo/KAGRA (LVK) experiment. The origin of these BBHs is unknown, and no electromagnetic (EM) counterpart has been undoubtedly associated to any of such GW events. The thin discs of active galactic nuclei (AGNs) might be viable environments where BBHs can form and at the same time produce observable radiation feedback. This paper presents new physically motivated light-curve (LC) solutions for thermal flares driven by the remnant of a BBH merger within the disc of an AGN. Following previous analyses, we consider that the BBH likely creates an under-density cavity in the disc prior to its coalescence. Depending on the merger conditions, the black hole (BH) remnant can leave the cavity, interact with the unperturbed disc, and drive a transient BH wind. The wind expels disc material that expands above and below the disc plane and we consider the emission of these plasma ejections as the EM counterpart to the merger GW. We model the LC of such eruptions as mini supernovae explosions finding that stellar mass merger remnants can drive distinguishable flares in optical and UV bands, with time lags of $\sim 10-100$ days after the GW event, when occurring in AGN with central engines of $\sim 10^{6-8}$ M. The Vera C. Rubin Observatory can potentially detect the optical component of these EM counterparts up to redshifts of $z\sim 0.7-1.2$, accordingly. The present model provide constraints on multi-messenger observables, such as the binary effective spin, mass ratio, remnant mass, and the time delay of the EM signature. These constraints are useful to localise flaring AGNs as sources of multi-messenger emission following GW alerts.
@@ I Introduction
Electromagnetic flares accompanying BBH mergers could unveil their astrophysical hosts. The search for such multi-messenger (MM) associations is the subject of current active research involving follow-up campaigns [8, 9, 10, 11, 12, 13], statistical analyses of observed data [11, 14, 15, 16, 17, 18], as well as theoretical efforts to understand the production of detectable electromagnetic counterparts [19, 20, 21, 22, 23, 24, 25, 26]. GWs and radiation from a common origin provide the means of measuring the distance and redshift of the source independently. In this regard, MM events triggered by BBH mergers can also find cosmological applications [27, 28, 29, 30]. Contrary to mergers of compact objects involving neutron stars, BBH mergers do not produce radiation by themselves and require occurring in gas-rich environments to produce significant EM signals.
The central regions of active galactic nuclei (AGNs) possess suitable conditions to cluster stellar mass black holes [31, 32, 33] and to induce BBH pairing and merging [34, 35, 36, 37, 38, 39, 40]. The scenario of mergers in AGN discs appears to naturally explain puzzling features revealed by LVK measurements, such as the anti-correlation of mass ratio and effective spin ($q-\chi_{\mathrm{eff}}$;  [41, 16]), as well as the production of intermediate mass black holes as merger remnants [42, 34]. Interestingly, AGN thin discs are environments dense enough where BBH mergers can produce significant radiation feedback [35, 21, 23].
@@
Figure 1: Sketch of the EM counterpart scenario discussed in this work. Upper: Radial cut of an AGN thin disc with vertical density stratification, where the BBH coalescence occurs within a pre-merger cavity. The merger remnant is kicked after coalescence and penetrates the unperturbed disc region. Middle: While travelling within the AGN disc, the remnant experiences a highly super-Eddington inflow of gas, leading to the production of a quasi-isotropic wind that expels the disc material above and below the disc plane. Lower: The plasma ejections produce an electromagnetic flare after expanding enough to allow the escape of internal photons.
This paper is organised as follows. In the next section, we derive an analytical model for the merger EM counterpart. In Section III, we present the spectra and optical light curves of the predicted emission and compare them with the sensitivity of optical instruments. We analyse constraints on the binary $q-\chi_{\mathrm{eff}}$ correlation and the flare time delay in Section IV. Finally, we provide a summary and conclusions in Section VI.
@@ II Electromagnetic counterpart: analytical considerations
Following such prediction and for analytical convenience, here we consider a BBH coalescence at the mid-plane of a thin disc, and within a cylindrical cavity of radius $r_{\mathrm{cav}}$, as depicted in Figure 1.
@@
A burst of GWs is released at coalescence, and the merger remnant is born with a mass $\sim M_{\bullet}$ [^1] The mass of the remnant is typically slightly smaller than the mass of the BBH due to the release of gravitational waves at coalescence. Nevertheless, for practical purposes proposed in this work, we do not make a distinction between the masses of the BBH and the remnant. with a likely gravitational kick velocity of magnitude $v_{\mathrm{k}}$ in the frame of the coalescence. The magnitude and direction of this gravitational recoil is sensitive to the mass ratio $q$ and spins of the binary components previous to the merger [57, 58, 59, 60], (we discuss the distribution of the kick magnitude and orientation in more detail in Section IV), and is generally highly supersonic with respect to the disc material.
Figure 2: Sketch illustrating the characteristic lengths and vertical heights $s_{i}$ defined in the text for calculating the disc mass expelled by the BBH merger remnant. Similar to the upper panel in Figure 1, the dashed thick line represents the trajectory of the merger remnant with angle $\theta_{\mathrm{k}}$ relative to the disc mid-plane. In the proposed scenario, the BH wind is ignited after the remnant travels the length $l_{\mathrm{HL}}=2GM_{\bullet}/c^{2}$ within the disc and is quenched upon reaching the height $s_{\mathrm{M,w}}=h_{\mathrm{d}}$. This transient wind expels the disc material located above and below the BH trajectory. The vertical dotted line indicates the wind quenching location along the BH trajectory and defines the volume of the ejected material in the proposed model (see the text).
@@
On the other hand, the radial drift velocity of the disc matter is highly subsonic [61], and thus the cavity crossing time of the remnant $r_{\mathrm{cav}}/v_{\mathrm{k}}$ is much shorter than the timescale of cavity replenishment [23]. Therefore, the interaction of the remnant with the unperturbed disc (and thus the possible production of electromagnetic feedback) crucially depends on the kick direction $\theta_{\mathrm{k}}$ with respect to the disc plane. We expect no interaction and hence no radiation counterpart for remnants with $\theta_{\mathrm{k}}\rightarrow\pi/2$. On the other hand, a significant interaction between the remnant and the unperturbed disc is expected for small $\theta_{\mathrm{k}}$ (see Figure 1), which is the case we focus on in this work.
@@
A kicked remnant of mass $M_{\bullet}$ penetrating the disc volume perturbs the disc material due to its gravitational influence. Such interaction can be initially assessed with Bondi-Hoyle-Lyttleton (BHL) theory. In a medium of gas density $\rho_{\mathrm{d}}$, the classical BHL inflow onto the kicked remnant is approximately $\dot{M}_{\mathrm{B}}\sim 4\pi\rho_{\mathrm{d}}(GM_{\bullet})^{2}/v_{\mathrm{k}}^{3}$. Normalising this inflow rate by the Eddington accretion limit of the remnant, gives for typical parameters of the problem:
@@
where $n_{\mathrm{d}}$ is the disc number density. Under such highly super-Eddington inflow, the remnant is expected to develop radiation-driven winds and/or bipolar collimated jets. Recent general relativistic magnetohydrodynamic (GRMHD) simulations of BHL accretion performed by [62] show the production of Blandford-Znajek jets [63] by spinning black holes (BHs) with a fixed spin direction, when traveling through a medium with plasma beta $\beta\gtrsim 10$. It is not clear, however, what the AGN disc’s $\beta$ is where BBHs likely coalesce. Furthermore, it is not clear whether the BBH coalescence produces a remnant with a fixed spin direction. BHs with precessing spin may launch precessing jets that could globally resemble wide-angle bipolar outflows. Given the aforementioned uncertainties for the production of powerful collimated jets, in the present work we conservatively consider that the remnant drives a wide-angle, bipolar, transient wind (i.e. a quasi-spherical outflow) during its trajectory within the disc. We assume that the remnant captures matter from the environment while simultaneously driving outflows. These inflow-outflow processes have been analysed in the context of dynamical friction of travelling black holes (BHs) with feedback [64, 65], and to explain the emission of ultra-luminous X-ray sources as supercritical accretors [66, 67].
The outcome of the remnant-disc interaction depends on the properties of the disc surrounding the merger, such as gas density, temperature, and disc thickness, which vary with the radial distance from the AGN central engine. One of the most employed models for describing AGN thin discs are those of Shakura & Sunyaev [61], Sirko & Goodman [68], and Thompson, Murray, & Quataert [69] (hereafter SS73, SG03, and TQM05, respectively). The SG03 model extends the pioneering work of SS73 by incorporating heating mechanisms to prevent material collapse due to gravitational instabilities at the outer disc regions. The TQM05 model address this by implementing mass loss due to star formation feedback (see [70] for a detailed comparison).
Considering the characteristics of the aforementioned disc models, in the present study we focus on AGN discs with relatively low accretion rates and following the SG03 description, for the following reasons. The SG03 model allows for a broader exploration of merger locations compared to SS73, whose validity is limited to regions where the disc remains gravitationally stable. While the TQM05 model provides a more accurate description for AGNs with high accretion rates, such systems are expected to produce substancial disc feedback. This feedback introduces additional sources of emission enhancement and flaring, complicating the detection of the EM counterparts discussed here. Finally, for AGNs with SMBHs ranging within $10^{6-8}$ M (including most SMBHs in the local Universe), the TQM05 model predicts significantly lower gas densities and smaller aspect ratios compared to the SG03 model, as shown in Appendix A. This suggests that the emission mechanism discussed here would produce weaker electromagnetic signatures in TQM05 discs than in SG03 discs.
At a given radial location $a$ in the disc, we consider the matter density as vertically stratified following a Gaussian profile (which is a suitable hydrostatic solution for the vertical disc structure):
$\rho_{\mathrm{d}}(a,s)=\rho_{\mathrm{c}}(a)\exp\left\{-\frac{1}{2}\left[\frac{s}{h_{\mathrm{d}}(a)}\right]^{2}\right\},$ (3)
being $\rho_{c}$ the density at the disc mid-plane and $h_{\mathrm{d}}$ the disc scale height, which depend on the radial location $a$.
In the present analysis, we asses the inflow onto the BH remnant based on the classical BHL theory. Therefore we assume this approach as valid when
$R_{\mathrm{HL}}=\frac{2GM_{\bullet}}{v_{\mathrm{k}}^{2}}<h_{\mathrm{d}}.$ (4)
@@
Once the remnant enters the disc, it takes a period to capture enough material to ignite the wind. We estimate this period as the crossing time of the BHL radius in the source frame:
$\Delta t^{\prime}_{\mathrm{HL}}=\frac{R_{\mathrm{HL}}}{v_{\mathrm{k}}}=\frac{2GM_{\bullet}}{v_{\mathrm{k}}^{3}}.$
Therefore, given the kick angle $\theta_{\mathrm{k}}$ with respect to the disc mid-plane, the BH wind is ignited when the remnant is located at the height:
$s_{\mathrm{m,w}}=\tan\theta_{\mathrm{k}}r_{\mathrm{cav}}+\sin\theta_{\mathrm{k}}R_{\mathrm{HL}}.$ (5)
above the mid-plane (see Figure 2). The wind is then quenched when the BH remnant captures insufficient inflow material to fuel the outflow. We approximate this quenching point when the remnant is located at the height:
$s_{\mathrm{M,w}}=h_{\mathrm{d}},$ (6)
above the mid-plane. At this height, the disc density has diminished to $60\%$ of its mid-plane value (see equation 3). At vertical locations $s>h_{\mathrm{d}}$, the BH can still experience inflow of disc material and produce outflows (although this feedback weakens as the disc gas density drops). Nevertheless, at $s>h_{\mathrm{d}}$, the disc atmosphere is strongly stratified (see equation 3), and the classical BHL theory (formulated for a massive body travelling within an external, uniform medium) does not apply. Given these conditions, we restrict our analysis to an outflow injected within the heights $s_{\mathrm{m,w}}$ and $s_{\mathrm{M,w}}$ (see Figure 2). Our analysis is then valid when $s_{\mathrm{m,w}}<s_{\mathrm{M,w}}$, which, using equations (5) and (6), results in the following condition:
$\tan\theta_{\mathrm{k}}\frac{r_{\mathrm{cav}}}{h_{\mathrm{d}}}+\sin\theta_{\mathrm{k}}\frac{R_{\mathrm{HL}}}{h_{\mathrm{d}}}<1.$ (7)
@@
The energy injected by the BH produces two expansive ejections of plasma above and below the accretion disc as depicted in Figure 1 (bottom panel). In this work, we consider the emission produced by such disc eruptions as the electromagnetic counterpart to the preceding merger gravitational waves (GWs), and we approximate the associated light curves (LCs) as those of explosive blast waves, similar to approaches in [71, 72, 25]. In Subsection II.1, we estimate the amount of matter and energy in the disc ejections, and in Subsection II.2 we assess their emission.
@@ II.1 The energy and mass of the disc ejections
During its passage through the disc, the remnant drives a transient outflow which we refer to as the ”BH wind”. This BH wind expels disc matter that emerges from the disc on the same and opposite sides relative to the direction where the remnant leaves the disc (see Figure 1, bottom). These ejections are referred to as the ”upper” and ”lower” eruptions, respectively.
@@
where $\rho_{\mathrm{d}}$ and $c_{\mathrm{s}}$ are the local density and speed of sound in the disc, which we obtain using the SG03 disc model.
@@
We assume equipartition of the total wind power for energising the upper and lower ejections of matter, which is a reasonable approximation for small kick angles $\theta_{\mathrm{k}}$. Thus, we calculate the energy stored by the BH wind in each ejection of plasma as: